# Trading Carbon for Physics: On the Resource Efficiency of Machine Learning for Spatio-Temporal Forecasting

### v1

@@ Abstract
Development of modern deep learning methods has been driven primarily by the push for improving model efficacy (accuracy metrics). This sole focus on efficacy has steered development of large-scale models that require massive resources, and results in considerable carbon footprint across the model life-cycle. In this work, we explore how physics inductive biases can offer useful trade-offs between model efficacy and model efficiency (compute, energy, and carbon). We study a variety of models for spatio-temporal forecasting, a task governed by physical laws and well-suited for exploring different levels of physics inductive bias. We show that embedding physics inductive biases into the model design can yield substantial efficiency gains while retaining or even improving efficacy for the tasks under consideration. In addition to using standard physics-informed spatio-temporal models, we demonstrate the usefulness of more recent models like flow matching as a general purpose method for spatio-temporal forecasting. Our experiments show that incorporating physics inductive biases offer a principled way to improve the efficiency and reduce the carbon footprint of machine learning models. We argue that model efficiency, along with model efficacy, should become a core consideration driving machine learning model development and deployment. [^1] Source code for the experiments in this work are available at: https://github.com/sophiawilson18/FlowMatching.
@@ 1 Introduction
Machine learning (ML) has advanced rapidly over the past decade (Sevilla et al., 2022), driven largely by the pursuit of improving model efficacy (accuracy metrics). This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle (Strubell et al., 2019; Anthony et al., 2020; Luccioni et al., 2023). As models scale, efficacy gains diminish while costs escalate, underscoring the need to move beyond accuracy as the sole measure. A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs (Henderson et al., 2022). This requires exploring trade-offs and identifying strategies that strike a favourable balance, rather than lopsided optimization.
@@
Figure 1: a: Conceptual map of modelling approaches by data availability (vertical axis) and domain knowledge (horizontal axis). Physics-informed ML lies in the middle, combining data-driven flexibility with physics priors. b: Test MSE versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles). The dashed line marks the Pareto front, dominated by the S-PINN, indicating the best accuracy-carbon trade-off. Details of the experiment are in Sec. 4.2.
Physics-informed ML offers a promising way to achieve this balance. Physics priors such as conservations laws, symmetries, or governing equations can be embedded into model design through, e.g., loss terms, architectural layers, or mechanisms. This combines the flexibility of data-driven methods with the structure of domain knowledge (Fig.1-a). As shown in Fig. 1-b, a semi-supervised physics-informed neural network (PINN) achieves the best accuracy–carbon trade-off, outperforming both a purely data-driven neural network (NN) and a physics-only PINN (see Sec. 4.1 for details). This example highlights how hybrid approaches can yield superior efficiency without sacrificing efficacy.
@@
Building on this intuition, we extend the perspective from task-specific models such as PINNs to general-purpose architectures that embed physics inductive biases in ways that are transferable across datasets and domains. We focus on spatio-temporal forecasting of partial differential equations (PDE) dynamics, a setting characterised by high computational cost and practical relevance. Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day (Lam et al., 2023). In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings, making this domain an ideal test-bed for exploring sustainable ML design (Van Wynsberghe, 2021). Our key contributions towards this end are:
@@
    Carbon-aware evaluation: We promote evaluation practices that consider both predictive accuracy and carbon cost, explicitly characterising trade-offs.
@@
    Characterising the accuracy–carbon trade-off: We conduct three experiments on PDE dynamics and spatio-temporal forecasting to compare a range of models with varying levels of physics inductive bias. Our main study focuses on forecasting incompressible shear flow, comparing U-nets (no/weak bias), flow matching (medium bias), and Fourier neural operators (strong bias).
Figure 2: Modelling the harmonic oscillator. Training and extrapolation predictions when using NNs with different activation functions: ReLU, Tanh, Sine, Snake, and using a PINN.
@@ 2 Background and Related Work
It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) (Kohavi et al., 1996). Striking the right balance is crucial to obtain the right class of models.
The same trade-off can also offer a useful way to perform model selection when viewed through the lens of resource consumption. Typically, if domain knowledge (physics-based or others) can be incorporated into models as inductive bias, it can reduce the reliance on training data or reduce prediction errors. This is illustrated in Fig. 2 for the task of modelling a harmonic oscillator. Models with appropriate inductive bias (sinusoidal oscillations) perform better for the same number of training data points and model parameters. See Sec. 4.1 for more details.
@@
The standard metrics used to characterise the performance of ML models primarily focus on efficacy and are untethered from efficiency considerations. This introduces a preference for models that could offer high efficacy at large resource costs (Bakhtiarifard et al., 2024). There are two primary streams of work that aim to include resource-aware metrics: using composite metrics and multi-objective optimisation.
Several composite metrics have been proposed to combine predictive performance with resources like energy consumption or carbon footprint. Evchenko et al. (2021) introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models. Carburacy (Moro et al., 2023) extends the idea to transformer models by jointly quantifying accuracy and carbon emissions. More recently, Kapoor et al. (2025) proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet (Tan and Le, 2019). More recently, energy consumption (Bakhtiarifard et al., 2024) and carbon footprint (Zhao et al., 2024) have been taken into consideration in NAS. These methods, however, have not actively studied the influence of physics inductive biases on the resource consumption of ML models.
@@
There are many model classes for the task of spatio-temporal forecasting; each differing in the extent of physics inductive bias they use. Fig. 3 provides a high-level overview of this span for models considered in this work.
Figure 3: Spectrum of model families from purely data-driven ML to physics-based solvers.
U-net. U-nets (Ronneberger et al., 2015) can be used for spatio-temporal forecasting by mapping a window of past states to future ones. They capture local correlations by fusing features across multiple spatial scales, but do not explicitly constrain the dynamics beyond what is enforced by the training data. Mild forms of inductive biases can be introduced to convolutional architectures like U-net by including circular padding to respect periodic boundaries and avoid edge discontinuities.
Flow Matching. Recent classes of probabilistic generative models like flow matching (FM) (Lipman et al., 2023) can be viewed as an intermediate class of models that can include some degree of inductive bias by modelling dynamics as continuous-time flows. This formulation resonates with dynamical systems and intuitively suggests smoother trajectories. FM has been applied to spatio-temporal forecasting and other PDE dynamics tasks (Lim et al., 2025; Li et al., 2025), with Baldan et al. (2025) further embedding PDE residuals into the objective.
Fourier Neural Operator. The Fourier Neural Operator (FNO) (Kovachki et al., 2023) introduces a physics prior through its Fourier-space representation, enabling the model to capture global spectral interactions. This makes it particularly well-suited to systems with periodicity and long-range correlations. In our spatio-temporal forecasting setting, we adapt the FNO from its original operator-learning formulation (Lu et al., 2021) to an autoregressive setup, where its spectral representation provides a strong inductive bias compared to the U-net by explicitly encoding global and periodic structure.
Physics-Informed Neural Network. PINNs (Raissi et al., 2019) embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency. By minimizing PDE residuals, PINNs can even be trained without supervised data, representing the strongest inductive bias among the models considered. Embedding conservation laws, symmetries, or governing PDEs into the ML model development pipeline has been shown to reduce the number of trainable parameters (Dutta et al., ; Patra et al., 2024), reduce the amount of training data (Psaros et al., 2023; Zhong et al., 2021), and accelerate convergence (Brehmer et al., 2024; Jahani-nasab and Bijarchi, 2024).
3 Methods
@@
We consider spatio-temporal forecasting of dynamical systems in the general setting, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given basic inputs such as initial and boundary conditions. We frame this as an autoregressive prediction task on a two-dimensional, uniformly spaced mesh $(x,y)\in\{x_{1},x_{2},\ldots,x_{H}\}\times\{y_{1},y_{2},\ldots,y_{W}\}$ with discretized time steps $t\in\{t_{1},t_{2},\ldots,t_{T}\}$. At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathbb{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$. $H$ and $W$ are spatial dimensions and $C$ is the number of physical fields (channels), such as pressure or velocity components.
To capture temporal dependencies, the model is conditioned on a finite history of length $h$. The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\ldots,u_{k}(x.y))$. A parameterised function $f_{\theta}$ is trained to predict:
@@
We adapt this generalised model formulation to three problems. Each problem is then tackled with relevant set of methods that range from being entirely data-driven to offering relevant degrees of physics inductive biases.
  1. 1.
    Harmonic Oscillator. Modelling the dynamics of the harmonic oscillator only requires the time variable i.e, $u(t)$. For the harmonic oscillator, Eq. 1 reduces to $u_{k+1}=f_{\theta}(u_{k-h+1:k}\big)$, where the spatial variables $(x,y)$ are dropped.
    We assess the impact of appropriate physics-inductive biases by comparing simple feed-forward NNs with and without biases. Specifically, these biases are introduced through activation functions and a loss term, for modelling the damped and undamped oscillators.
  2. 2.
    Viscous Burgers’ Equation. Solving the viscous Burgers’ equation increases the modelling complexity compared to the harmonic oscillator, as it has one spatial and one temporal variable to be considered. Eq. 1 for this case is given as: $u_{k+1}(x)=f_{\theta}(u_{k-h+1:k}(x)\big)$, where only one spatial variable $x$ is modelled along time.
    We evaluate how different combinations of data and physics constraints influence predictive accuracy and efficiency by comparing a supervised NN, a physics-only PINN, and a hybrid PINN for approximating the viscous Burgers’ equation.
  3. 3.
    Incompressible Shear Flow. Modelling the spatio-temporal dynamics of incompressible shear flow governed by Navier-Stokes equations is the most complex task considered. It has dependence on both the spatial and temporal variables. As a result, we use Eq. 1 in its original formulation.
    We evaluate several models spanning a broad spectrum of physics biases, depicted in Fig. 3. Purely data-driven models like U-net, the moderately biased model FM, and strongly biased neural operators are adapted for this task. The U-nets and FNOs use a history of $h=4$ states, i.e. they are trained for one-step prediction according to Eq. 1 i.e., $u_{k+1}(x,y)=f_{\theta}(u_{k-3:k}(x,y)\big).$
    FM replaces direct one-step prediction with a continuous transport between consecutive states. Instead of learning a map $u_{k-h+1:k}(x,y)\mapsto u_{k+1}(x,y)$, FM predicts the next state by transporting the current state forward along a learned flow. Starting from $u_{k}(x,y)$, the method defines a flow trajectory $\phi_{\tau}(u_{k})$ over a synthetic time axis $\tau\in[0,1]$, where the flow evolves under a time-dependent vector field $v_{\tau}$. The one-step forecast is then obtained as the endpoint of this flow: $\hat{u}_{k+1}(x,y)=\phi_{1}(u_{k}(x,y))$, where $\partial_{\tau}\phi_{\tau}(u)=v_{\tau}(\phi_{\tau}(u)),\phi_{0}(u)=u_{k}(x,y).$ In practice, a NN is trained to approximate this vector field, defined by an interpolation path between consecutive states. Further FM details including the interpolation path, vector field, and learning objective are provided in App. A.
4 Data and Experiments
Based on the problem formulation in Sec. 3, we set-up experiments on the three tasks to explore the trade-off between efficacy and efficiency. Specific choices for each of the experiments are elaborated in detail next.
4.1 Harmonic Oscillator
We use simple feed-forward NNs to model the harmonic oscillator by introducing physics biases in two ways: through activation functions, where Sine and Snake (Ziyin et al., 2020) have periodic biases, and through a physics-informed loss objective. ReLU and Tanh serve as unbiased baselines.
The training data is generated from the damped harmonic oscillator, $u(t)=e^{-\delta t}(2A\cos(\phi+\omega t))$, where: $A=1/(2\cos\phi),\penalty 10000\ \penalty 10000\ \phi=\arctan\left(-\delta/\omega\right)\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \omega=\sqrt{\omega_{0}^{2}-\delta^{2}}\penalty 10000\$, with added Gaussian noise $(\mu=0,\sigma=0.3)$. We consider both the non-damped ($\delta=0$) and the damped ($\delta=0.02$) regimes.
All models use the same underlying architecture (two hidden layers of width 64), trained for $100$ epochs with Adam optimizer (Kingma and Ba, 2015) using a batch size of 25. Learning rates are $0.01$ for standard networks and $0.001$ for the PINN, which employs Snake activations. Results are averaged over $10$ runs. Additional details are provided in App. B.
4.2 Viscous Burgers’ Equation
Figure 4: Target and training data for viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residuals samples. d: Initial and boundary condition points.
In this experiment, we study models to approximate the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$
$u(x,0)=-\sin(\pi x),$
$u(1,t)=u(-1,t)=0.$
for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$). The domain is discretized with $256$ spatial points and $100$ time steps. We compare three models with varying physics inductive biases, all using the same underlying architecture (comprising four hidden layers, width $50$, tanh activation functions):
  • NN: a supervised NN trained with the mean-squared error (MSE) on ground-truth data, using the loss objective: $\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}$.
  • U-PINN: an unsupervised PINN trained only on physics constraints that minimizes PDE residuals, initial conditions (IC) and boundary conditions (BC). Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial/boundary conditions, expressed as: $\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{PDE}}+\mathcal{L}_{\text{IC}}+\mathcal{L}_{\text{BC}}$.
  • S-PINN: a semi-supervised PINN where half of the training points are used for PDE residuals and boundary/initial conditions, while the other half enforce a data-driven penalty (MSE). Its loss objective is given as, $\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{U-PINN}}+\mathcal{L}_{\text{NN}}$.
The specific formulations of all three loss objectives are provided in App. C. All models were trained using Adam optimizer followed by L-BFGS optimizer (Liu and Nocedal, 1989) with learning rate $10^{-3}$. Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN. Data and code are adapted from Khoa (2022).
4.3 Incompressible Shear Flow
We set-up the large-scale experiment on modelling incompressible shear flow governed by the Navier–Stokes equations, using the spatio-temporal forecasting models introduced in Section 2.
@@ 4.3 Incompressible Shear Flow
The incompreesible shear flow data is derived from The Well dataset (Ohana et al., 2024) generated with Dedalus (Burns et al., 2020). We use a subset of the full dataset consisting of 240 simulations spanning six PDE parameter settings with $40$ initial conditions each. Simulations are resolved at $256\times 512$ with $200$ time steps ($\Delta t=0.1$). The data is split by initial condition into 80/10/10 for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout (initialised at $t=0$). The dataset is particularly challenging due to the coupled evolution of four physical fields (velocity components, pressure, tracer) and its non-linear spatiotemporal dynamics at high resolution. Further details on PDE formulation and initialization are provided in App. D.
Model(s) Physics inductive bias Strength
@@
Table 1: Overview of the eight models and the relative strength of their physics inductive biases.
@@
We evaluate eight architectures spanning a spectrum of physics inductive biases, from purely data-driven to strongly physics-informed, summarised in Table 1. We use standard U-nets (Ronneberger et al., 2015) and a ConvNeXt variant (UNet-CN) (Liu et al., 2022) for improved hardware-efficiency. Zero-padding serves as our no-bias baseline, while circular padding (CP) introduces a weak prior consistent with periodic boundary conditions in the dataset. We also include our proposed FM model. Finally, we include three variants of the Fourier Neural Operator (FNO): a baseline model, a Tucker-factorized (TFNO) variant (Kossaifi et al., 2023) that improves parameter effiency, and a U-shaped (UFNO) variant (Rahman et al., 2023) designed to better capture multi-scale interactions.
@@
Across all models, we use a batch size of 16 and the AdamW optimizer (Loshchilov and Hutter, 2019) with $10^{-4}$ weight decay. For U-nets and FNOs, we adopt the coarse-tuned learning rates reported in The Well ($5\cdot 10^{-4}$ and $10^{-3}$, respectively). For FM, we performed a separate coarse tuning and selected $5\cdot 10^{-3}$. A linear warm-up cosine scheduler was applied for the first three epochs. The loss optimised is MSE averaged across fields and space. Training ran on a single Nvidia A40 GPU ($\leq$ 24h) with early stopping (patience 6). FM was trained in low-resolution data space with subsequent upsampling, avoiding the additional overhead of an autoencoder. Further details on FM in low-resolution data space are provided in App. E, and complete model configurations are listed in App. F.
4.4 Energy and Carbon Metrics
We employ Carbontracker (Anthony et al., 2020) to estimate the energy consumption and carbon footprint during training and inference by monitoring real-time hardware power usage (CPU, GPU, DRAM) and adjusting for infrastructure overhead via the power usage effectiveness. The carbon footprint is reported as carbon dioxide equivalent (CO2eq) emissions, which converts all greenhouse gases into an equivalent CO2 warming potential. It is derived from the total energy consumed and by using the global average carbon intensity for 2024 at $445$ gCO2/kWh, ensuring fair comparisons across models (IEA, 2025).
Figure 5: Damped harmonic oscillator predictions. Mean predictions (solid lines) with standard deviation (shaded areas) from $10$ runs. Results are shown for the two best-performing models: Snake and PINN.

### v2

# Physics Priors Offer Useful Accuracy-Carbon Trade-Offs in Spatio-Temporal Forecasting

@@ Abstract
Development of modern deep learning methods has been driven primarily by the push for improving model efficacy (accuracy metrics). This sole focus on efficacy has steered development of large-scale models that require massive computational resources, and results in considerable energy consumption and corresponding carbon footprint across the model lifecycle. In this work, we explore how physics inductive biases can offer useful trade-offs between model efficacy and model efficiency (compute, energy, and carbon). We study models with strong, weak, and no physics-inductive biases for spatio-temporal forecasting of incompressible shear flow, a task governed by the Navier-Stokes equations. We find that models with stronger physics priors achieve substantially lower training footprints, but this advantage does not straightforwardly extend to inference, highlighting the importance of evaluating carbon costs across the full model lifecycle rather than any single stage. We argue that model efficiency, along with model efficacy, should become a core consideration driving machine learning model development and deployment.
@@ 1 Introduction
Machine learning (ML) has advanced rapidly over the past decade, driven largely by the pursuit of improving model efficacy (accuracy metrics). This narrow focus has led to increasingly large models that require substantial computational resources and generate significant carbon emissions across their lifecycle [1, 2, 3, 4]. As models scale, efficacy gains diminish while costs escalate, underscoring the need to move beyond accuracy as the sole measure. A more holistic assessment must balance efficacy with efficiency, capturing not only predictive performance but also compute, energy, and carbon costs [5]. This requires identifying strategies that strike a favourable balance rather than optimising for accuracy alone.
@@
Figure 1: a: Conceptual map of modelling approaches by data availability (vertical axis) and domain knowledge (horizontal axis). Physics-informed ML lies in the middle, combining data-driven flexibility with physics priors. b: Test mean squared error (MSE) versus CO2eq emissions for a data-driven NN (red squares), an unsupervised PINN (blue dots), and a semi-supervised PINN (grey triangles). The dashed line marks the Pareto front, dominated by the S-PINN, indicating the best accuracy-carbon trade-off. Details of the experiment are in App. 0.A.
Physics-informed ML offers a promising path towards this balance. By embedding prior knowledge about the underlying system directly into the model, one can combine the flexibility of data-driven methods with the structure of domain knowledge, as illustrated in Fig. 1-a. Prior work has shown that embedding conservation laws, symmetries, or governing equations can reduce the number of trainable parameters [6, 7], reduce the amount of training data [8, 9, 10], and accelerate training convergence [11, 12].
Whether physics-informed ML models can also reduce energy consumption and carbon footprint remains a less explored area. Fig. 1-b offers a concrete illustration of its potential. In a simple experiment on a partial differential equation (PDE) system, we compare three models that share the same underlying architecture but differ in how much they rely on physics through their loss functions: a purely data-driven neural network (NN), an unsupervised physics-informed neural network (PINN) trained solely on PDE residuals, and a semi-supervised PINN combining data and physics. The semi-supervised PINN dominates the Pareto front, achieving lower prediction error at a lower carbon cost than either alternative. This suggests that striking the right balance between data and physics does not just improve accuracy, but can simultaneously reduce the carbon footprint. Full details of this experiment are provided in App. 0.A.
@@
Building on this intuition, we extend the perspective from task-specific models such as PINNs to general-purpose architectures that embed physics inductive biases in ways that are transferable across datasets and domains. We focus on spatio-temporal forecasting of PDE dynamics, a setting characterised by high computational cost and practical relevance. Operational weather and climate forecasting, for instance, requires large-scale PDE solvers running on supercomputers multiple times per day [13]. In such contexts, even modest efficiency improvements can translate into substantial operational carbon savings [14], making this domain an ideal test-bed for exploring sustainable ML design [15]. Our key contributions towards this end are:
@@
    Carbon-aware evaluation: We promote evaluation practices that explicitly account for training and inference costs alongside predictive accuracy.
@@
    Characterising the accuracy-carbon trade-off: We provide an empirical characterisation of how varying degrees of physics priors affect the accuracy-carbon trade-off in spatio-temporal forecasting, evaluated on a complex, high-resolution dataset of incompressible shear flow.
@@ 2 Background and Related Work
Figure 2: Bias-variance trade-off. Top: classical trade-off with minimum total error at optimal model complexity. Bottom: possible analogous balance between physics-based constraints and data-driven flexibility.
@@
It is commonly understood in ML that strong model assumptions (bias) risk under-fitting, whereas high sensitivity to training data can lead to over-fitting (variance) [16]. Striking the right balance is crucial to obtain the right class of models.
We draw an analogy to model design under resource constraints. Incorporating domain knowledge as inductive bias can reduce reliance on large datasets and potentially lower computational demands, but may limit model flexibility. Conversely, highly data-driven models offer flexibility, but must learn all patterns from data which often requires lots of data and computational resources. We therefore hypothesise that a physics-data balance may exist, analogous to the bias-variance trade-off (Fig. 2).
@@
The standard metrics used to characterise the performance of ML models primarily focus on efficacy and are untethered from efficiency considerations. This introduces a preference for models that could offer high efficacy at large resource costs [17]. There are two primary streams of work that aim to include resource-aware metrics: using composite metrics and multi-objective optimisation.
Several composite metrics have been proposed to combine predictive performance with resources like energy consumption or carbon footprint. Evchenko et al. [18] introduced a composite metric that combines efficacy and efficiency metrics to study the influence of resource constraints on traditional (non deep learning) ML models. Carburacy [19] extends the idea to transformer models by jointly quantifying accuracy and carbon emissions. More recently, Kapoor et al. [20] proposed EcoL2, a composite metric similar to Carburacy but extended to include the costs of data generation and hyperparameter tuning.
Multi-objective optimisation has been primarily used in neural architecture search (NAS) resulting in models like EfficientNet [21]. More recently, energy consumption [17] and carbon footprint [22] have been taken into consideration in NAS. These methods, however, have not actively studied the influence of physics inductive biases on the resource consumption of ML models.
@@
There are many model classes for the task of spatio-temporal forecasting, differing in the extent of physics inductive bias they use. In this work, we focus on two: the U-net [23] and the Fourier Neural Operator (FNO) [24].
U-nets perform forecasting by mapping a window of past states to future states. They capture local correlations by fusing features across multiple spatial scales, but do not impose any explicit constraints on the underlying dynamics beyond those learning from data. Mild forms of inductive biases can be introduced to convolutional architectures like U-net, for example through circular padding to respect periodic boundaries and avoid edge discontinuities.
The FNO incorporates a stronger inductive bias through its Fourier-space representation, enabling the model to capture global spectral interactions. This makes it particularly well-suited to systems with periodicity and long-range correlations. In our setting, we adapt the FNO from its original operator-learning formulation [25] to an autoregressive setup, where its spectral parameterisation explicitly encodes global and periodic structure, in contrast to the primarily local bias of the U-net.
3 Experiment
@@
We consider spatio-temporal forecasting of dynamical systems, where the objective is to approximate the solution $u(x,y,t)$ of a PDE given initial and boundary conditions. We frame this as an autoregressive prediction task on a two-dimensional, uniformly spaced mesh $(x,y)\in\{x_{1},x_{2},\ldots,x_{H}\}\times\{y_{1},y_{2},\ldots,y_{W}\}$ with discretized time steps $t\in\{t_{1},t_{2},\ldots,t_{T}\}$. At each step $t_{k}$ the state is represented as a tensor $u_{k}(x,y)\in\mathrm{R}^{H\times W\times C}$, where $k\in\{1,2,..,T\}$ indexes discrete time steps in a trajectory of length $T$. $H$ and $W$ are spatial dimensions and $C$ is the number of physical fields (channels), such as pressure or velocity components.
To capture temporal dependencies, the model is conditioned on a finite history of length $h$. The input window is denoted as $u_{k-h+1:k}(x,y):=(u_{k-h+1}(x,y),\\ \ldots,u_{k}(x.y))$. A parameterised function $f_{\theta}$ is trained to predict:
@@
We instantiate this framework for forecasting incompressible shear flow governed by the Navier-Stokes equations, a setting that is both computationally demanding and governed by well-understood physical structure, making it well-suited for studying the efficacy-efficiency trade-off.
@@
We use the incompressible shear flow data from The Well collection [26], generated with Dedalus [27]. The dataset captures the coupled non-linear evolution of four physical fields: horizontal and vertical velocity, pressure, and a passive tracer on a periodic spatial domain, see example in Fig. 3. We work with a subset, which consists of 240 simulations spanning six PDE parameter settings and $40$ initial conditions settings. Simulations are resolved at $256\times 512$ with $200$ time steps ($\Delta t=0.1$). The data is split by initial condition into 8080/1010/$10$ for training, validation, and testing, yielding $37,632$ one-step samples for training, $4,704$ for one-step inference, and $24$ trajectories for autoregressive rollout initialised at $t=0$. Further details on the PDE formulation and initialisation are provided in App. 0.B.
Table 1: Overview of the seven models and the relative strength of their physics inductive biases.
Models Physics inductive bias Strength
@@
We evaluate seven models spanning a spectrum of physics inductive biases, summarised in Table 1. We consider a baseline U-net and a ConvNeXt variant (UNet-CN) [28], which improves hardware efficiency through modern convolutional design. We introduce a weak form of inductive bias by applying circular padding to both architectures (UNet-CP, UNet-CN-CP). For models with stronger physics priors, we include three FNO variants: a baseline FNO, a Tucker-factorized variant (TFNO) [29] that improves parameter efficiency, and a U-shaped variant (UFNO) [30] designed to better capture multi-scale interactions.
@@
All models use a history of $h=4$ states and are trained for one-step prediction according to Eq. 1. Across all models, we use a batch size of 16 and the AdamW optimizer [31] with $10^{-4}$ weight decay. We adopt the coarse-tuned learning rates reported in The Well [26] ($5\cdot 10^{-4}$ for U-nets and $10^{-3}$ for FNOs, respectively), which were tuned across several datasets including the one used here. A linear warm-up cosine scheduler was applied for the first three epochs and the loss is the mean squared error (MSE) averaged across fields and space. Training ran on a single Nvidia A40 GPU ($\leq$ 24h) with early stopping (patience 6). Full model configurations are listed in App. 0.C.
Carbon Footprint.
To estimate energy consumption during training and inference, we employ Carbontracker [2]. We report results in terms of carbon footprint rather than energy consumption directly, since energy is only problematic to the extent that it is generated from carbon-emitting sources. Were the grid entirely renewable, the same energy expenditure would carry negligible direct carbon emissions. In practice, however, the global energy mix remains far from carbon-neutral, and we therefore convert energy to carbon dioxide equivalent (CO2eq) emissions using the global average carbon intensity for 2024 at $445$ gCO2/kWh [32]. Using a fixed carbon intensity ensures fair comparisons across models, independent of when or where they were trained.